OpenStudy (anonymous):

Find the maximum of the object function f, subject to the restraints. (i have writte out the problem in a reply)

OpenStudy (anonymous):

\[f=15x+25y\] Restraints \[3x+4y \ge 60\] \[x+8y \ge 40\] \[11x+28y \le 380\] \[x \ge 0, y \ge0\]

OpenStudy (dumbcow):

graph the restraints as inequalities and find the "shaded region" so to speak max will be at one of the intersection points of "shaded region"